OpenStudy (anonymous):

Help! Can anyone help me prove this i tried so hard! 1 - 3 cos(x) - 4 cos²(x) = (1 + cos(x))(1 - 4 cos(x)) sin²(x) = 1 - cos²(x) = (1 + cos(x))(1 - cos(x))

OpenStudy (anonymous):

Call x = cos x, the first trinomial is in the form: -4x^2 - 3x + 1. It can be factored as: (x + 1)(x - 1/4). TIP to do it. This is Tip 1 in solving quadratic equation. When a - b + c = 0, one root is (-1) and the other is (-c/a). Replace x by cos x to get the first identity. The second identity is very popular: (1 - x^2) = (1- x)(1 + x).

OpenStudy (anonymous):

Correction: The trinomial can be factored as (x + 1)(1/4 - x) because a is -4.